Factor each number or monomial completely.
51
-63
41
-95

Answer by stanbon(48558)

Factor each number or monomial completely.
51 = 3*17
-63 = -3^2*7
41 = 41 is prime
-95 = -5*19
===============
Cheers,
Stan H.